# Cache Coherence Over Disaggregated Memory

### v3

@@ Abstract.
Disaggregating memory from compute offers the opportunity to better utilize stranded memory in cloud data centers. It is important to cache data in the compute nodes and maintain cache coherence across multiple compute nodes. However, the limited computing power on disaggregated memory servers makes traditional cache coherence protocols suboptimal, particularly in the case of stranded memory. This paper introduces SELCC; a Shared-Exclusive Latch Cache Coherence protocol that maintains cache coherence without imposing any computational burden on the remote memory side. It aligns the state machine of the shared-exclusive latch protocol with the MSI protocol by introducing lazy latch-release and invalidation messages, thereby ensuring both atomicity of data access and cache coherence. SELCC embeds cache-ownership metadata directly into the RDMA latch word, enabling efficient cache ownership management via RDMA atomic operations. SELCC can serve as an abstraction layer over disaggregated memory with APIs that resemble main-memory accesses. A concurrent B-tree and three transaction concurrency control algorithms are realized using SELCC’s abstraction layer. Experimental results show that SELCC significantly outperforms Remote-Procedure-Call-based protocols for cache coherence under limited remote computing power. Applications on SELCC achieve comparable or superior performance over disaggregated memory compared to competitors.
@@ 1. Introduction
Memory disaggregation has emerged as a significant trend in cloud databases in both academia (Zhang et al., 2020a, b; Korolija et al., 2022; Wang and Zhang, 2023; Lu et al., 2024; Zuo et al., 2021; Luo et al., 2023; Wang et al., 2023a) and industry (Cao et al., 2021; Zhang et al., 2021; Yang et al., 2024). An important motivation behind disaggregated memory is to utilize the substantial amounts of stranded memory (Zhang et al., 2022; Gu et al., 2017; Rao and Porter, 2016; Grandl et al., 2014; Verma et al., 2015) in cloud data centers. Stranded memory refers to memory that is inaccessible due to all the available cores being allocated to virtual machines (Zhang et al., 2022). Memory disaggregation addresses this issue by accessing the stranded memory via high-speed networks, physically decoupling the memory resources from compute servers. By establishing disaggregated memory over stranded memory, cloud providers can significantly enhance memory utilization and reduce the total cost of ownership (TCO).
@@
Our Approach. This paper presents the Shared-Exclusive Latch-based Cache-Coherence protocol (SELCC), a new protocol for the cache coherence problem over disaggregated memory. By introducing lazy latch-release and invalidation messages, the one-sided RDMA shared-exclusive latch protocol can be upgraded to address the cache-coherence problem with sequential consistency. It embeds the cache ownership information into RDMA latch words, allowing both the latch and cache ownership to be managed within a single RDMA atomic operation. To optimize performance, SELCC prioritizes local ownership handover over global ownership handover, employs batched processing for cache eviction, and minimizes the bandwidth for dirty data flush back. Additionally, SELCC enhances the fairness by incorporating priority into invalidation messages. In Fig. 1c, SELCC operates without involving remote memory computing power, making it particularly suited for memory disaggregation. Note that in scenarios where remote memory has computing resources, they can be used for other functionalities (rather than handling cache coherence), such as operator pushdown, which can result in significant performance gains (Lu et al., 2024; Wang et al., 2023b).
@@ 3. System Overview
This paper addresses optimizing DBMSs for disaggregated memory (Wang et al., 2023a; Yang et al., 2024; Li et al., 2024), where compute and memory are decoupled, and multiple compute nodes share a common memory pool (Fig. 2).
@@ 4.3.1. Embeded Ownership in RDMA Latch Word
In SELCC, we propose to embed cache ownership within the RDMA latch words, allowing both cache ownership and RDMA latch to be managed via one RDMA atomic operation. As in Fig. 5, a 64-bit latch word is the maximum data length supported by RDMA atomic operations. We divide these 64 bits into 2 parts: (1) An exclusive latch holder’s ID (8 bits), and (2) A reader holders’ ID bitmap (56 bits). [^1] This protocol can support up to 56 compute nodes. With multi-cores on each compute node, a system using SELCC can support thousands of cores.
@@ 4.3.2. Acquiring Ownership via One-Sided RDMA
Acquiring “Shared” Ownership. To acquire shared ownership, a reader atomically fetches the latch word and sets its own position in the bitmap using RDMA_FAA (with the add operand set to $(1<<NodeID)$). The reader checks the return value of RDMA_FAA to determine if a writer is currently holding the latch. If so, the acquisition fails, and the reader resets its bit in the bitmap using another RDMA_FAA. Then. it sends an invalidation message based on the holder ID of the exclusive latch.
Upgrading “Shared” to “Modified” Ownership. When a writer finds that the target GCL is cached in “Shared” state, the compute node needs to upgrade the global latch from shared to exclusive. First, the compute node attempts to compare and swap the global latch words from $(0,1<<NodeID)$ to $(NodeID,00..0)$. If two nodes simultaneously upgrade the same global latch, both nodes will continuously fail at the CAS operation. This issue is analogous to the deadlock problem for lock upgrade within a single machine, and is resolved using a typical fallback approach. [^2] Note that global lock upgrading is always triggered implicitly, and the atomic ownership/latch upgrade for the user is not necessary. After several failed attempts to upgrade the latch, SELCC resorts to releasing the shared latch followed and acquiring the exclusive latch.
@@ 4.4.1. Case 1: A Writer on the Sender Side Invalidates a “Modified” Copy on the Receiver Side
Global Ownership Handover allows the receiver of an invalidation message to transfer ownership directly to the requester through one atomic RDMA_FAA operation, rather than requiring the sender to repeatedly attempt latch acquisition (Fig. 7b). For example, if Node $A$ (Writer) issues an invalidation message to Node $B$ (“Modified” state), where $A$ and $B$ are the Node IDs, ownership can be atomically transferred from Node $B$ to Node $A$ by adding the value $(A-B,00..0)$ onto the latch word via RDMA_FAA.
@@ 4.4.2. Case 2: A Reader on the Sender Side Invalidates a “Modified” Copy on the Receiver Side
Refer to Fig. 7 for illustration. Node $A$ attempts to read a GCL while Node $B$ holds a copy in the “Modified” state. This scenario is generally handled in a manner similar to Case 1, with minor adjustments for global ownership handover and GCL forwarding. As in Fig. 7b, when the current owner in the “Modified” state receives an invalidation message, it atomically updates the latch word by adding $(-B,(1<<A)|(1<<B))$ via RDMA_FAA. This operation clears the modified ownership for Node $B$ and assigns shared ownership to both Nodes $A$ and $B$. Afterwards, the latest GCL is forwarded to the sender, and it is also flushed back to the disaggregated memory (unlike Case 1). This ensures that future concurrent readers can directly acquire a “Shared” copy from the disaggregated memory rather than receiving the GCL forwarded from another compute node.
@@ 5.1. Efficient Invalidation Messages
Invalidation messages, implemented via RPC over RDMA, play a critical role in coordinating global cache-line (GCL) ownership in SELCC. Invalidation messages contain key information, e.g., the global address of the target GCL and the case of invalidation (Sec. 4.4). As in Fig. 9, the RPC request is sent via RDMA_send, to the receiver side maintaining a ring buffer to capture incoming messages via RDMA receive. The background message handlers in the receiver node process the invalidation messages by releasing/handovering the global latch. After processing, the handler sends an acknowledgment along with the latest GCL copy to a local buffer on the sender’s side via RDMA write. Importantly, RPC only exists between compute servers, while communication between compute and memory servers remains strictly one-sided.
@@ 5.1.1. Lower Priority for Processing Invalidation Messages
In scenarios where invalidation messages are handled concurrently with front-end accessing threads, synchronization is required for message handlers to manage access to cache frame metadata. A straightforward approach is to acquire the local exclusive latch before processing the invalidation messages. However, this can lead to two problems: (1) the message handler can get blocked if the local accessing thread holds the local latch for an extended period, preventing the processing of subsequent invalidation messages, and (2) such a design assigns equal priority to local accessing threads and global accessing threads for handovering the data ownership. In workloads with high contention, as illustrated in Fig. 10 left, this can lead to frequent ownership transfers among compute nodes, significantly increasing read/write latency and generating a large volume of invalidation messages.
@@ 5.2.1. Batched Processing for GCL Eviction
To address these issues, we implement a ring buffer for each queue pair connected to remote memory. The ring buffer size is set to match the maximum number of outstanding RDMA work requests per queue pair, $L_{out}$. Before executing RDMA operations, the eviction worker attempts to acquire an available slot from the ring buffer and copies the payload from the cache frame into this buffer slot. If no slot is available, the worker must busy-wait for the completion of previous RDMA operations. Upon receiving $n$ completion notifications from the completion queue, the ring buffer’s tail is advanced by $n$ positions. This design ensures that the number of outstanding RDMA requests always be within the threshold $L_{out}$, and the cache frames can be immediately recycled once the dirty GCL is copied to the ring buffer.
@@ 5.3. Fairness
Fairness is a significant challenge for the SELCC protocol, as it is based on a shared-exclusive spinlock. The read and write latency on a particular server can become extremely long if that server experiences starvation during latch acquisition. Next, we present the root causes of latch starvation over SELCC and propose relevant solutions accordingly. Due to the two-level hierarchy of the system, two root causes of latch starvation can be identified, each requiring distinct resolution techniques.
@@ 5.3.1. Handling Local Latch Starvation
To address local latch starvation, we implement a lease mechanism that forces the compute node to release the global latch when a data object has been continuously accessed by local front-end threads for an extended period. To interrupt these continuous local accesses at an appropriate time, two counters, the read access counter ($R_{c}$) and the write access counter ($W_{c}$), are maintained in each cache entry. These counters are activated only when an invalidation message is dropped due to the ongoing local access and is deactivated when a thread acquires the latch without spinning, indicating that the data is no longer heavily accessed. The counters are incremented by 1 when a local access waits for the latch. Synthetic access times for the cache entry are calculated by $A_{times}=\frac{R_{c}}{P}+W_{c}$, where $P$ represents the number of front-end threads on the compute node. When the synthetic access times $A_{times}$ exceed a predefined threshold $\theta$, the local thread proactively release the global latch and reset the counters.
@@ 5.3.2. Handling Global Latch Starvation
To handle global latch starvation, we adopt a priority aging mechanism, originally devised to solve the starvation problem in CPU scheduling (Tanenbaum, 2009). In SELCC, each invalidation message is assigned a priority that is positively correlated to the number of retries a compute node has conducted for a particular RDMA latch. The idea is to ensure that global latch ownership is handed over to the compute node with the highest latch priority. To achieve this, we implement 3 mechanisms.
First, as in Sec. 5.1, there is a manually injected time interval between each retry for a particular latch. This interval decreases as the priority of latch acquisition increases. Thus, compute nodes with prolonged wait times are more likely to successfully acquire the latch through more frequent latch retries. Second, the global ownership handover mechanism introduced in Sec. 4.4.1 takes starvation priority as a key factor in determining the next owner of the data. The exclusive latch holder, receiving invalidation messages from all conflicted servers, acts as a centralized decision-maker for global latch ownership transfer. During the continuous local access (Sec. 5.3.1), the invalidation message handler keeps receiving invalidation messages from other compute nodes and memorizes those messages as well as their priority. Upon handover the global ownership, it is deliberately transferred to the sender with the highest priority. Finally, to address write starvation caused by continuous global reads, all global readers create a time window during which no concurrent reader holds the target shared latch, allowing the concurrent writer to acquire the exclusive latch. To realize this, when latch starvation is detected, a flag is set in the cached frame. This flag triggers a spinning wait before the next global latch acquisition for the GCL. The spin duration is calculated as $T_{spin}=P_{inv}\times RTT$, where $P_{inv}$ is the priority level of the received invalidation message and $RTT$ is the round trip time for RDMA atomic operations.
@@ 6.1. Programming Interface
SELCC exposes a straightforward interface to upper-level applications (Table 1). Users can allocate or deallocate global cache lines by calling Allocate/Free. Each data access is conducted via the local cache, and has to be protected by an SELCC latch that consists of a hierarchical data structure containing a local latch in the cache frame and a global latch in the remote memory. The acquisition of a SELCC latch (SELCC_SLock/SELCC_XLock) ensures that both the local and global latch are obtained, thereby guaranteeing access atomicity and cache coherence across compute nodes. Upon acquisition, the API returns a cache handle pointing to the local copy of the target GCL. The release of the SELCC latch (SELCC_SUnLock/SELCC_XUnLock) ensures the immediate release of the local latch while deferring the release of the global latch until another compute node accesses the same GCL. Additionally, SELCC provides APIs for global atomic operations that can be utilized to generate global timestamps or sequential numbers.

### v4

@@ Abstract.
Disaggregating memory from compute offers the opportunity to better utilize stranded memory in cloud data centers. It is important to cache data in the compute nodes and maintain cache coherence across multiple compute nodes. However, the limited computing power on disaggregated memory servers makes traditional cache coherence protocols suboptimal, particularly in the case of stranded memory. This paper introduces SELCC; a Shared-Exclusive Latch Cache Coherence protocol that maintains cache coherence without imposing any computational burden on the remote memory side. It aligns the state machine of the shared-exclusive latch protocol with the MSI protocol , thereby ensuring both atomicity of data access and cache coherence with sequential consistency. SELCC embeds cache-ownership metadata directly into the RDMA latch word, enabling efficient cache ownership management via RDMA atomic operations. SELCC can serve as an abstraction layer over disaggregated memory with APIs that resemble main-memory accesses. A concurrent B-tree and three transaction concurrency control algorithms are realized using SELCC’s abstraction layer. Experimental results show that SELCC significantly outperforms Remote-Procedure-Call-based protocols for cache coherence under limited remote computing power. Applications on SELCC achieve comparable or superior performance over disaggregated memory compared to competitors.
@@ 1. Introduction
Memory disaggregation has emerged as a significant trend in cloud databases in both academia (Zhang et al., 2020a, b; Korolija et al., 2022; Wang and Zhang, 2023; Lu et al., 2024; Zuo et al., 2021; Luo et al., 2023; Wang et al., 2023b) and industry (Cao et al., 2021; Zhang et al., 2021; Yang et al., 2024). An important motivation behind disaggregated memory is to utilize the substantial amounts of stranded memory (Zhang et al., 2022; Gu et al., 2017; Rao and Porter, 2016; Grandl et al., 2014; Verma et al., 2015) in cloud data centers. Stranded memory refers to memory that is inaccessible due to all the available cores being allocated to virtual machines (Zhang et al., 2022). Memory disaggregation addresses this issue by accessing the stranded memory via high-speed networks, physically decoupling the memory resources from compute servers. By establishing disaggregated memory over stranded memory, cloud providers can significantly enhance memory utilization and reduce the total cost of ownership (TCO).
@@
Our Approach. This paper presents the Shared-Exclusive Latch-based Cache-Coherence protocol (SELCC), a new protocol for the cache coherence problem over disaggregated memory. By introducing lazy latch-release and invalidation messages, the one-sided RDMA shared-exclusive latch protocol can be upgraded to address the cache-coherence problem with sequential consistency. It embeds the cache ownership information into RDMA latch words, allowing both the latch and cache ownership to be managed within a single RDMA atomic operation. To optimize performance, SELCC prioritizes local ownership handover over global ownership handover, employs batched processing for cache eviction, and minimizes the bandwidth for dirty data flush back. Additionally, SELCC enhances the fairness by incorporating priority into invalidation messages. In Fig. 1c, SELCC operates without involving remote memory computing power, making it particularly suited for memory disaggregation. Note that in scenarios where remote memory has computing resources, they can be used for other functionalities (rather than handling cache coherence), such as operator pushdown, which can result in significant performance gains (Lu et al., 2024; Wang et al., 2023c).
@@ 3. System Overview
This paper addresses optimizing DBMSs for disaggregated memory (Wang et al., 2023b; Yang et al., 2024; Li et al., 2024), where compute and memory are decoupled, and multiple compute nodes share a common memory pool (Fig. 2).
@@ 4.3.1. Embeded Ownership in RDMA Latch Word
In SELCC, we propose to embed cache ownership within the RDMA latch words, allowing both cache ownership and RDMA latch to be managed via one RDMA atomic operation. As in Fig. 5, a 64-bit latch word is the maximum data length supported by RDMA atomic operations. We divide these 64 bits into 2 parts: (1) An exclusive latch holder’s ID (6 bits), and (2) A reader holders’ ID bitmap (58 bits). [^1] This protocol can support up to 58 compute nodes. With multi-cores on each compute node, a system using SELCC can support thousands of cores.
@@ 4.3.2. Acquiring Ownership via One-Sided RDMA
Acquiring “Shared” Ownership. To acquire shared ownership, a reader atomically fetches the latch word and sets its own position in the bitmap using RDMA_FAA (with the add operand set to $(1<<NodeID)$). The reader checks the return value of RDMA_FAA to determine if a writer is currently holding the latch. If so, the acquisition fails, and the reader resets its bit in the bitmap using another RDMA_FAA. Then it sends an invalidation message based on the holder ID of the exclusive latch.
Upgrading “Shared” to “Modified” Ownership. When a writer finds that the target GCL is cached in “Shared” state, the compute node needs to upgrade the global latch from shared to exclusive. First, the compute node attempts to compare and swap the global latch words from $(0,1<<NodeID)$ to $(NodeID,00..0)$. If two nodes simultaneously upgrade the same global latch, both nodes will continuously fail at the CAS operation. This issue is analogous to the deadlock problem for lock upgrade within a single machine, and is resolved using a typical fallback approach. [^2] Note that global lock upgrading is always triggered implicitly, and the atomic ownership/latch upgrade for the user is not necessary. After several failed attempts to upgrade the latch, SELCC resorts to releasing the shared latch and then acquiring the exclusive latch.
@@ 4.4.1. Case 1: A Writer on the Sender Side Invalidates a “Modified” Copy on the Receiver Side
Global Ownership Handover allows the receiver of an invalidation message to transfer ownership directly to the requester through one atomic RDMA_FAA operation, rather than requiring the sender to repeatedly attempt latch acquisition (Fig. 7b) (Johnson et al., 2009; Joshi, 1991). For example, if Node $A$ (Writer) issues an invalidation message to Node $B$ (“Modified” state), where $A$ and $B$ are the Node IDs, ownership can be atomically transferred from Node $B$ to Node $A$ by adding the value $(A-B,00..0)$ onto the latch word via RDMA_FAA.
@@ 4.4.2. Case 2: A Reader on the Sender Side Invalidates a “Modified” Copy on the Receiver Side
Refer to Fig. 7 for illustration. Node $A$ attempts to read a GCL while Node $B$ holds a copy in the “Modified” state. This scenario is generally handled in a manner similar to Case 1, with minor adjustments for global ownership handover and GCL forwarding. As in Fig. 7b, when the current owner in the “Modified” state receives an invalidation message, it atomically modifies the latch word by adding $(-B,(1<<A)|(1<<B))$ via RDMA_FAA. This operation clears the modified ownership for Node $B$ and assigns shared ownership to both Nodes $A$ and $B$. Additionally, in the same RDMA round trip, the latest GCL is flushed back to the disaggregated memory together (unlike Case 1), ensuring that future concurrent readers can acquire the latest “Shared” copy from the disaggregated memory rather than fetch it from another compute node. Finally, the latest GCL is forwarded to the sender.
@@ 5.1. Efficient Invalidation Messages
Invalidation messages, implemented via RPC over RDMA, play a critical role in coordinating global cache-line (GCL) ownership in SELCC. Invalidation messages contain key information, e.g., the global address of the target GCL and the case of invalidation (Sec. 4.4). As in Fig. 9, the RPC request is sent via RDMA_send, to the receiver side maintaining a ring buffer to capture incoming messages via RDMA receive. The background message handlers in the receiver node process the invalidation messages by releasing/handing over the global latch. After processing, the handler sends an acknowledgment along with the latest GCL copy to a local buffer on the sender’s side via RDMA write. Importantly, in SELCC, RPC is used only between compute nodes, while the communication between compute and memory layers are strictly one-sided. SELCC adheres to the compute-free design principle for memory nodes, distinguishing it fundamentally from RPC-based protocols.
@@ 5.1.1. Lower Priority for Processing Invalidation Messages
In scenarios where invalidation messages are handled concurrently with front-end accessing threads, synchronization is required for message handlers to manage access to cache frame metadata. A straightforward approach is to acquire the local exclusive latch before processing the invalidation messages. However, this can lead to two problems: (1) the message handler can get blocked if the local accessing thread holds the local latch for an extended period, preventing the processing of subsequent invalidation messages, and (2) such a design assigns equal priority to local accessing threads and global accessing threads for handing over the data ownership. In workloads with high contention, as illustrated in Fig. 10 left, this can lead to frequent ownership transfers among compute nodes, significantly increasing read/write latency and generating a large volume of invalidation messages.
@@ 5.2.1. Batched Processing for GCL Eviction
To address these issues, we implement a ring buffer for each queue pair connected to remote memory. The ring buffer size is set to match the maximum number of outstanding RDMA work requests per queue pair, $L_{out}$. Before executing RDMA operations, the eviction worker attempts to acquire an available slot from the ring buffer and copies the payload from the cache frame into this buffer slot. If no slot is available, the worker must busy-wait for the completion of previous RDMA operations. Upon receiving $n$ completion notifications from the completion queue, the ring buffer’s tail is advanced by $n$ positions. This design ensures that the number of outstanding RDMA requests always be within the threshold $L_{out}$, and the cache frames can be immediately recycled once the dirty GCL is copied to the ring buffer. However, these performance gains are not achieved without cost. The introduced ring buffer incurs a space overhead that scales linearly with both the number of threads per node and the number of compute nodes in the system.
@@ 5.3. Fairness
Fairness is a significant challenge for the SELCC protocol, as it is based on a shared-exclusive spinlock. A server may experience starvation if it fails to acquire the latch repeatedly. Next, we present the root causes of latch starvation over SELCC and propose relevant solutions accordingly. Due to the two-level hierarchy of the system, two root causes of latch starvation can be identified, each requiring distinct resolution techniques.
@@ 5.3.1. Handling Local Latch Starvation
To address local latch starvation, we implement a lease mechanism that forces the compute node to release the global latch when a data object has been continuously accessed by local front-end threads for an extended period. To interrupt these continuous local accesses at an appropriate time, two counters, the read access counter ($R_{c}$) and the write access counter ($W_{c}$), are maintained in each cache entry. These counters are activated only when an invalidation message is dropped due to the ongoing local access and is deactivated when a thread acquires the latch without spinning, indicating that the data is no longer heavily accessed. The counters are incremented by 1 when a local access waits for the latch. Synthetic access times for the cache entry are calculated by $A_{times}=\frac{R_{c}}{P}+W_{c}$, where $P$ represents the number of front-end threads on the compute node. When the synthetic access times $A_{times}$ exceed a predefined threshold $\gamma$, the local thread proactively release the global latch and reset the counters.
@@ 5.3.2. Handling Global Latch Starvation
To handle global latch starvation for asymmetric hardware among compute nodes, we adopt a priority aging mechanism, originally devised to solve the starvation problem in CPU scheduling (Tanenbaum, 2009). In SELCC, each invalidation message is assigned a priority that is positively correlated to the number of retries a compute node has conducted for a particular RDMA latch. The global ownership handover mechanism introduced in Sec. 4.4.1 takes starvation priority as a key factor in determining the next owner of the data. The exclusive latch holder, receiving invalidation messages from all conflicted servers, acts as a centralized decision-maker for global latch ownership transfer. During the continuous local access (Sec. 5.3.1), the invalidation message handler keeps receiving invalidation messages from other compute nodes and memorizes those messages as well as their priority. Upon handing over the global ownership, it is deliberately transferred to the sender with the highest priority. Furthermore, as in Sec. 5.1, there is a manually injected time interval between each retry for a particular latch. This interval decreases as the priority of latch acquisition increases. Thus, compute nodes with prolonged wait times are more likely to successfully acquire the latch through more frequent latch retries.
To addressed global write starvation induced by continuous global reads, the protocol employs a spin-waiting mechanism. Specifically, when a high-priority invalidation message is detected on reader nodes, a flag is set in the corresponding local cached frame, compelling subsequent global readers to spin for a predetermined duration. The spin duration of the global readers corresponds to the priority of the starved writer. The spin duration is proportional to the priority of the starved writer, thereby creating a sufficiently large time window during which no concurrent reader holds the targeted shared latch and enabling a concurrent writer to preempt latch ownership. Although this approach addresses starvation, readers still achieve higher throughput owing to their inherently greater concurrency relative to writers. To balance performance between read and write operations, when write starvation is detected, its priority is recorded in the cache frame; consequently, this GCL cache frame must be invalidated by a global reader with equal or greater priority. By this priority matching mechanism, the performance of global readers and writers can be balanced given skewed workloads (See Fig. 15 b) Although the measurements above address the fairness issue, they also introduce additional space overhead, approximately 70 bytes per cached GCL.
@@ 6.1. Programming Interface
SELCC exposes a straightforward interface to upper-level applications (Table 1). Users can allocate or deallocate global cache lines by calling Allocate/Free. Each data access is conducted via the local cache, and has to be protected by an SELCC latch that consists of a hierarchical data structure containing a local latch in the cache frame and a global latch in the remote memory. The acquisition of a SELCC latch (SELCC_SLock/SELCC_XLock) ensures that both the local and global latch are obtained, thereby guaranteeing access atomicity and cache coherence across compute nodes. Upon acquisition, the API returns a cache handle pointing to the local copy of the target GCL. Due to the lazy latch-release introduced in Sec. 4.1, the release of the SELCC latch (SELCC_SUnLock/SELCC_XUnLock) only ensures the immediate release of the local latch while deferring the release of the global latch until another compute node accesses the same GCL. Additionally, SELCC provides APIs for global atomic operations that can be utilized to generate global timestamps or sequential numbers.